# VideoLifter: Lifting Videos to 3D with Fast Hierarchical Stereo Alignment

### v1

@@ Abstract
Efficiently reconstructing accurate 3D models from monocular video is a key challenge in computer vision, critical for advancing applications in virtual reality, robotics, and scene understanding. Existing approaches typically require pre-computed camera parameters and frame-by-frame reconstruction pipelines, which are prone to error accumulation and entail significant computational overhead. To address these limitations, we introduce VideoLifter, a novel framework that leverages geometric priors from a learnable model to incrementally optimize a globally sparse to dense 3D representation directly from video sequences. VideoLifter segments the video sequence into local windows, where it matches and registers frames, constructs consistent fragments, and aligns them hierarchically to produce a unified 3D model. By tracking and propagating sparse point correspondences across frames and fragments, VideoLifter incrementally refines camera poses and 3D structure, minimizing reprojection error for improved accuracy and robustness. This approach significantly accelerates the reconstruction process, reducing training time by over 82% while surpassing current state-of-the-art methods in visual fidelity and computational efficiency.
@@ 1 Introduction
Efficiently reconstructing 3D scenes from sequential image observations is a longstanding challenge in computer vision, with applications spanning virtual and augmented reality (VR/AR), video processing, outdoor photogrammetry, autonomous driving, urban planning, and medical imaging. Recently, reconstructing 3D scenes from single video sequences has gained significant traction. This trend is driven by two factors: the increasing accessibility of handheld capture devices, which makes video capture more practical for non-professional users, and recent advancements in high-fidelity 3D reconstruction techniques from images. Notable examples include Neural Radiance Fields (NeRF) [26] and 3D Gaussian Splatting [18], which have expanded the possibilities for capturing complex, high-quality 3D representations from image data.
@@
Figure 1: Novel View Synthesis and Training Time Comparisons. VideoLifter does not require precomputed camera parameters (i.e., camera intrinsics K from COLMAP), reduces the training time required by the most relevant baseline CF-3DGS by 82% while improving image quality (SSIM) by 0.2.
@@
Most video-to-3D reconstruction methods, either based on radiance fields or Gaussian splitting, heavily depend on Structure-from-Motion (SfM) [32] to generate initial sparse reconstructions, providing essential components like camera poses, intrinsics, and an initial point cloud to build dense 3D models. However, in video data, SfM is not reliable or even approachable (Issue ➀) due to its reliance on photometric assumptions that often fail in low-texture or challenging lighting conditions [24, 27, 32].
To address the limitations of traditional SfM, recent methods [12, 6, 21, 42] have shifted toward jointly optimizing camera poses and scene representations rather than relying on initial sparse reconstructions from SfM. These methods generally show a working manner of incrementally reconstructing scenes from a canonical view. However, they face several significant limitations: (a) Lack of Robustness (Issue ➁): they still rely on accurate camera intrinsics from SfM, which makes them unreliable when applied to in-the-wild video data. (b) Slow and Inefficient (Issue ➂): these methods are very slow due to their iterative, frame-by-frame approach requiring re-optimizing the entire sequence with each new frame. Moreover, for camera pose optimization, they assume constant-velocity camera motion and also re-optimize camera poses from scratch, which prolongs training times and complicates handling complex trajectories. (c) Susceptibility to Incremental Errors (Issue ➃): in long video sequences, the frame-by-frame approach makes these methods prone to cumulative errors.
To address the limitations of existing methods, we propose VideoLifter, a novel approach for efficient and robust video-to-3D reconstruction that operates without the need for SfM. Our method follows the typical sparse-to-dense strategy with two key innovations: (1) Sparse point-based fragment registration: inspired by hierarchical SfM [34], we capture a sparse scene representation from video data by dividing the input video into non-overlapping fragments anchored by key frames. We construct an efficient subgraph from these key frames to optimize anchor point clouds, providing shared camera intrinsics and transformation matrices across fragments. Local camera extrinsics within each fragment are estimated through sparse feature matching, leveraging pretrained geometric priors from foundation model (e.g. MASt3R [20]) and eliminating the need for precomputed camera poses or intrinsic from unreliable SfM (Issue ➀ ➁). (2) Hierarchical 3D Gaussian alignment: for dense 3D reconstruction, we address cumulative errors (Issue ➃) over long video sequences by introducing a hierarchical approach. Starting with the registered fragments, local 3D Gaussians are initialized and optimized based on photometric consistency. These local representations are then hierarchically merged into a unified scene representation, ensuring global consistency and accurate camera pose alignment.
By leveraging our fragment-by-fragment and hierarchical strategy, VideoLifter achieves both high computational efficiency and high quality renderings when handling video sequences (Issue ➂). Moreover, the lack of reliance on precomputed camera parameters or SfM makes our approach highly robust and flexible. As shown in Fig. 1, VideoLifter achieves >5×>5$\times$> 5 × speed improvements and enhanced view-synthesis quality compared to various state-of-the-art methods.
@@
  • We introduce VideoLifter, an efficient, flexible, and robust video-to-3D reconstruction framework, reinventing the sparse-to-dense reconstruction strategy with two novel stages.
  • Our sparse point-based fragment registration efficiently captures the sparse representation for long-sequence videos in a fragment-by-fragment manner, utilizing a learnable model for geometric priors without relying on SfM, ensuring robustness across video data.
  • Our hierarchical 3D Gaussian alignment minimizes incremental errors while significantly enhancing reconstruction speed, achieving more than 5× speed improvements over existing methods.
  • Extensive experiments on the Tanks and Temples and CO3D-V2 datasets demonstrate that VideoLifter significantly improves training efficiency and rendering quality compared to state-of-the-art methods.
2 Related Works and Preliminary
@@
3D reconstruction for high-quality novel view synthesis is the task of generating unseen views of a scene or object from a set of given images [1, 25]. One of the most influential methods, Neural Radiance Fields (NeRF) [26], introduced the use of multilayer perceptrons (MLPs) to represent 3D scenes. By taking the 3D coordinates and viewing directions as input and utilizing volume rendering techniques, NeRF achieves photo-realistic image synthesis. However, despite its ability to produce high-quality renderings, NeRF is computationally expensive and slow, both during training and inference. Research following NeRF has aimed at either improving the quality of renderings [2, 3, 4] or boosting the efficiency of the approach [30, 33, 37, 36], though few methods succeed in addressing both quality and efficiency simultaneously. This is largely due to the inherently ill-posed nature of transforming a sparse point cloud into a dense 3D scene representation. In recent years, unstructured radiance field methods have emerged [7, 18, 44], using different scene representation primitives. Among these, 3D Gaussian Splatting (3D-GS)[18] stands out by employing anisotropic 3D Gaussians[46] to model radiance fields. This technique, combined with differentiable splatting for rendering, has demonstrated impressive performance in efficiently reconstructing complex, real-world scenes with high fidelity. However, 3D-GS performs the sparse-to-dense optimization starting from structure-from-motion point clouds, named adaptive density control, which is critical for converting sparse point clouds into dense 3D Gaussian representations. A notable limitation is that the optimization of SfM and the subsequent 3D Gaussian refinement are decoupled, which can lead to the accumulation of errors that the current pipelines are unable to correct. NeRFs and 3DGS rely on carefully captured sequential video or multi-view images to ensure sufficient scene coverage, utilizing preprocessing tools like Structure-from-Motion (SfM) software such as COLMAP [32] to compute camera parameters and provide a sparse SfM point cloud as additional input.
@@
Estimating 3D structure and camera motion simultaneously is a well-explored challenge, with a long-established set of techniques addressing the problem [13, 29, 43]. Structure from Motion (SfM) has seen significant advancements across various dimensions. Methods like [22, 15] focus on enhancing feature detection, while [35] introduces innovative optimization approaches. Improved data representations and more robust structural solutions are explored by [14, 32]. Simultaneous Localization and Mapping (SLAM) systems, on the other hand, track camera motion and construct 3D structures from video sequences [29, 9, 28]. In [9], the photometric bundle adjustment algorithm is proposed, which directly minimizes photometric errors by aligning pixels. Despite these advances, traditional SfM and SLAM techniques remain vulnerable to issues such as low-texture regions, occlusions, moving objects, and lighting variations, limiting their overall robustness and performance.
@@
The limitations of SfM—particularly in handling textureless surfaces, motion blur, or specular reflections—can lead to cumulative errors that propagate through subsequent stages of radiance field reconstruction, reducing overall accuracy. To overcome these issues, several approaches have been proposed to eliminate the reliance on SfM during the optimization of radiance fields. NeRFmm [42], for example, jointly optimizes camera intrinsics, extrinsics, and radiance field training, bypassing the need for pre-computed poses. Similarly, BARF [21] adopts a coarse-to-fine pose refinement strategy while jointly optimizing NeRF to handle pose inaccuracies. SCNeRF [16] incorporates a parameterization of camera distortions and utilizes geometric loss to regularize ray tracing. GARF [16] further simplifies joint pose and scene optimization by leveraging Gaussian-MLP models to improve both efficiency and accuracy. SPARF [39] and TrackNeRF [23] introduces a method to simulate pose noise by injecting Gaussian noise into the camera parameters. More recent work integrates depth information as an additional constraint to refine radiance field and 3DGS optimization. Nope-NeRF [6], Lu-NeRF [8], LocalRF [24], and CF-3DGS [12] all utilize depth data to guide the optimization process. Pushing the envelope further, new techniques such as the work in [17] and InstantSplat [10] employ monocular depth estimators and image-matching networks, and integrate end-to-end stereo models like DUST3R [40] to reduce the dependency on camera pose information. While these methods show promise in reducing the number of views required for accurate scene reconstruction, scaling them effectively to a large number of views remains a challenge.
3D Gaussian Splatting
3D Gaussian Splatting [18] is an explicit 3D scene representation utilizing a set of 3D Gaussians to model the scene. A 3D Gaussian is parameterized by a mean vector $\bm{x}\in{}^{3}$ and a covariance matrix $\Sigma\in{}^{3\times 3}$:
$\small G(\bm{p})=\frac{1}{\left(2\pi\right)^{3/2}\left|\Sigma\right|^{1/2}}e^{-\frac{1}{2}{\left(\bm{p}-\bm{x}\right)}^{T}\Sigma^{-1}\left(\bm{p}-\bm{\mu}\right)}$ (1)
To represent the view-direction-dependent, spherical harmonic (SH) coefficients are attached to each Gaussian, and the color is computed via $\bm{c}\left(\bm{d}\right)=\sum_{i=1}^{n}{\bm{c}}_{i}\mathcal{B}_{i}\left(\bm{d}\right)$, where $\mathcal{B}_{i}$ is the $i^{\rm{th}}$ SH basis. And the color is rendered via $\bm{c}=\sum_{i=1}^{n}\bm{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$, where $\bm{c}_{i}$ is the color computed from the SH coefficients of the $i^{\rm{th}}$ Gaussian. $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance multiplied by a learned per-Gaussian opacity. The 2D covariance matrix is calculated by projecting the 3D covariance to the camera coordinates. The 3D covariance matrix is decomposed into a scaling matrix and a rotation matrix and is optimized as Gaussian attributes.
@@ 3 VideoLifter: An Efficient and Effective Video-to-3D Framework
We first start with the definition of the targeted video-to-3D reconstruction problem that VideoLifter solves.
Problem
Given a sequence of $N$ unposed and uncalibrated images from a monocular video, denoted as =$\mathcal{I}=\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$, VideoLifter reconstructs the scene using 3D Gaussians $\mathcal{G}$ along with estimated camera intrinsics $K$ and extrinsics $\mathcal{T}=\{T_{i}\in\mathbb{R}^{3\times 4}\}_{i=1}^{N}$. Here, we assume all frames share a common intrinsic matrix, as a single monocular video capture is the input.
Challenges
Existing frameworks often rely on a frame-by-frame approach, making them inherently slow and prone to cumulative errors when processing videos. Furthermore, these methods depend on Structure-from-Motion for camera intrinsics, which can be unreliable or even infeasible for in-the-wild video data due to low texture, inconsistent lighting, or motion blur.
Figure 2: Network Architecture. Our method takes uncalibrated images as input and reconstructs a dense scene representation based on self-supervised photometric signals. It performs efficient sparse reconstruction through the use of “feature points” to formulate an initial solution for globally aligned camera poses. It subsequently optimizes a set of 3D Gaussians and corrects erroneous estimations from each fragment, hierarchically aligning them into a globally coherent 3D representation.
Our High-level Framework
To ensure both efficiency and effectiveness while handling video data, we propose a novel sparse-to-dense reconstruction pipeline, as illustrated in Fig. 2. The pipeline integrates two key innovations designed for efficiency and robustness:
  • Sparse Point-Based Fragment Registration with 3D Priors enables efficient sparse reconstruction in a fragment-by-fragment way and also eliminates reliance on SfM by using 3D priors.
  • Hierarchical 3D Gaussian Alignment obtains a globally coherent 3D representation in a hierarchical way to both boost efficiency as well as reduce merging error.
@@ 3.1 Fragment Registration with Learned 3D Priors
To enable efficient sparse reconstruction, we propose a fragment-by-fragment approach, diverging from conventional methods that process video either frame-by-frame or by optimizing the entire video sequence simultaneously. This design choice significantly enhances the computational efficiency of our pipeline. Specifically, we divide the input sequence into $m$ disjoint windows of length $k$ (referred to as fragments). For example, the $i$’th fragment would be $\mathcal{I}^{f}_{i}=[I_{(i-1)k+1},I_{(i-1)k+2},\dots,I_{ik}]$, where $i\in[1,m]$.
To process data in a fragment-by-fragment manner, it is essential to extract both intra-fragment and inter-fragment informative features. For intra-fragment processing, we estimate the locally aligned camera extrinsics $\mathcal{T}^{f}_{i}$, within fragment $\mathcal{I}^{f}_{i}$, using the sparse point cloud:
$\small\mathcal{T}^{f}_{i}=\text{Frag-Regis}(\mathcal{I}^{f}_{i},f_{\text{prior}}(\mathcal{I}^{f}_{i})),\quad i\in[1,m],$
where $f_{\text{prior}}$ refers to the SfM which is widely used to initialize reconstruction. However, due to the inefficiency and unreliability of SfM in video settings, we propose leveraging a large-scale 3D model to obtain robust 3D geometric priors [20].
Furthermore, to ensure the consistency of point clouds between adjacent fragments, we “anchor” each fragment using its first frame, $I_{(i-1)k+1}$, referred to as the key frame. Key frames are used to obtain a globally consistent optimized dense point map, as well as intrinsic $K$ and transformation matrices across adjacent fragments $\{T^{f}_{i\rightarrow i+1}\}_{i=1}^{m-1}$.
Model Inter-Fragment Consistency with Efficient Key Frame Optimization
Processing the video in a fragment-wise manner, rather than frame-by-frame, requires careful handling of inter-fragment consistency, since each fragment independently estimates its extrinsics without direct inference from other fragments.
For a monocular video, a straightforward solution to build inter-fragment consistency is to build a complete graph by taking every pair as the input of MASt3R [20] to enforce global consistency. However, rather than constructing a complete graph, we build a more efficient sub-graph by only using key frames and the four closest frames to each key frame. This approach leverages the fact that MASt3R’s stereo predictions are more accurate for pairs with significant co-visibility. Specifically, given a set of key frames $\{I_{(i-1)k+1}\}_{i=1}^{m}$, where $k$ is the number of views per fragment and $m$ is the number of fragments, we construct a subgraph $\mathcal{G}_{m}(\mathcal{V},\mathcal{E})$ to model the relationships among key frames and their neighboring frames. In this subgraph, each vertex $v\in\mathcal{V}$ corresponds to a key frame or one of its four closest neighboring frames, and each edge $e=(v,u)\in\mathcal{E}$ represents shared visual content between frames $I_{v}$ and $I_{u}$.
Key frame optimization is then performed to achieve three objectives: optimized point maps for key frames $\tilde{\bm{P}}$, relative transformations $\bm{T}_{e}$, and a shared focal $\bar{f}$. Here $\{T^{f}_{i\rightarrow i+1}\}_{i=1}^{m-1}\subset\bm{T}_{e}$.
The optimization of the transformation matrix, scaling factor $\sigma_{e}$, and point map) is carried on the efficient sub-graph. Following the procedure in MASt3R, for each image pair $e=(v,u)\in\mathcal{E}$, we denote the point maps $\bm{P}_{v,v}$ and $\bm{P}_{u,v}$, along with their confidence maps $\bm{O}_{v,v}$ and $\bm{O}_{u,v}$. For clarity, we define $\bm{P}_{v,e}:=\bm{P}_{v,v}$, $\bm{P}_{u,e}:=\bm{P}_{u,u}$, $\bm{O}_{v,e}:=\bm{O}_{v,v}$, and $\bm{O}_{u,e}:=\bm{O}_{u,u}$. To this end, we optimize
@@
Note that this efficient sub-graph could be extended to long video sequences without excessive computational demands, as the complexity of this sub-graph is bounded by $\mathcal{O}(\frac{N}{k}*4)$, which scales linearly with the number of images $N$.
With the optimized point map, the focal length for each camera could be further computed following MASt3R:
$\small f^{*}=\arg\min_{f}\sum_{i=0}^{W}\sum_{j=0}^{H}\bm{O}^{i,j}\left\|(i^{\prime},j^{\prime})-f\frac{(\bm{P}^{i,j,0},\bm{P}^{i,j,1})}{\bm{P}^{i,j,2}}\right\|.$
Assuming a single-camera setup for monocular video, similar to COLMAP for single-scene capture, we propose estimating the shared focal length by averaging across all training views: $\bar{f}=\frac{1}{N}\sum_{i=1}^{N}f_{i}^{*}$.
Feature Tracking within Fragment
The input video is partitioned into non-overlapping fragments, each containing $k$ frames. Within each fragment, pairwise feature correspondences are established between the key frame and all subsequent frames, reducing the incremental errors commonly encountered in sequential matching.
@@
To refine the camera poses derived from initial pairwise estimates, consider the first fragment $\mathcal{I}^{f}_{1}=\{I_{1},I_{2},\dots,I_{k}\}$ as an example. We begin by identifying the intersection of 2D correspondences between the key frame $I_{1}$ and each subsequent frame from index 2 to $k$. This process yields a consistent set of correspondences visible across all frames within the same fragment. Using these intersected 2D correspondences, the corresponding 3D positions from the key frame, previously optimized during key frame processing, are retrieved. These 3D-2D correspondences serve as input to PnP-RANSAC [11], refining the camera poses to ensure alignment with consistent 3D points across views in the fragment.

### v2

@@ Abstract
Efficiently reconstructing 3D scenes from monocular video remains a core challenge in computer vision, vital for applications in virtual reality, robotics, and scene understanding. Recently, frame-by-frame progressive reconstruction without camera poses is commonly adopted, incurring high computational overhead and compounding errors when scaling to longer videos. To overcome these issues, we introduce VideoLifter, a novel video-to-3D pipeline that leverages a local-to-global strategy on a fragment basis, achieving both extreme efficiency and SOTA quality. Locally, VideoLifter leverages learnable 3D priors to register fragments, extracting essential information for subsequent 3D Gaussian initialization with enforced inter-fragment consistency and optimized efficiency. Globally, it employs a tree-based hierarchical merging method with key frame guidance for inter-fragment alignment, pairwise merging with Gaussian point pruning, and subsequent joint optimization to ensure global consistency while efficiently mitigating cumulative errors. This approach significantly accelerates the reconstruction process, reducing training time by over 82% while holding better visual quality than current SOTA methods.
@@ 1 Introduction
Efficiently reconstructing 3D scenes from sequential image observations is a longstanding challenge in computer vision, with applications spanning AR/VR, video processing, and autonomous driving. Recently, reconstructing 3D scenes from single video sequences (video-to-3D) has gained significant traction. This trend is driven by two factors: the increasing accessibility of handheld capture devices, making video capture more practical for non-professional users, and recent advancements in high-fidelity 3D reconstruction methods such as Neural Radiance Fields (NeRF) [30] and 3D Gaussian Splatting (3D-GS) [21].
@@
Figure 1: Novel View Synthesis and Training Time Comparisons. VideoLifter does not require precomputed camera parameters (i.e., camera intrinsics K from COLMAP), reduces the training time required by the most relevant baseline CF-3DGS [14] by 82% while improving image quality (SSIM).
@@
Most video-to-3D reconstruction methods, based on either NeRF or 3D-GS, heavily depend on Structure-from-Motion (SfM) [36] to generate initial sparse reconstructions, providing essential components like camera poses, intrinsics, and an initial point cloud to build dense 3D models. However, when applied to video data, SfM is often unreliable or even infeasible (Issue ➊), because it relies on photometric assumptions that frequently break down in low-texture or challenging lighting conditions [28, 31, 36], though some works improve SfM for some specific conditions [25]. In response, recent methods [14, 6, 24, 46] have shifted toward jointly optimizing camera poses and scene representations rather than relying solely on SfM-based initializations. Nevertheless, these approaches still depend on accurate camera intrinsics from SfM, limiting their applicability in in-the-wild video scenarios.
More importantly, those SfM-free video-to-3D methods typically reconstruct scenes incrementally from a canonical view, with two critical issues. First, they are slow and inefficient (Issue ➋) due to an iterative, frame-by-frame approach that re-optimizes the entire sequence with each new frame, thereby prolonging training times ($>$ 2 hours) and complicating the handling of complex trajectories, especially given the video setup, not a few images. Second, they are susceptible to incremental errors (Issue ➌), as the frame-by-frame approach tends to accumulate errors over long video sequences.
To address these issues, we propose VideoLifter, a novel video-to-3D reconstruction pipeline that achieves a 5$\times$ speed-up and enhanced view-synthesis quality compared to state-of-the-art methods, as demonstrated in Fig. 1.
We effectively adopt the local-to-global stream to handle long-sequence videos on a fragment basis and then subsequently merge fragments into a final, globally consistent 3D scene. Our pipeline is driven by two key innovations that make the local-to-global concept workable with significantly boosted efficiency (Issue ➋) and much-reduced incremental errors(Issue ➌). First, in the Fragment Registration with Learned 3D Priors (Local) stage, we extract essential information (e.g., pointmaps and local camera poses for 3D Gaussian initialization) from each fragment by leveraging pretrained prior models such as MASt3R [23] to address Issue ➊. Rather than naively using 3D priors to initialize 3D Gaussians, like InstantSplat [12], we improve efficiency by (1) enforcing inter-fragment consistency via solely considering the key frames (set to first frame), solved on an efficient subgraph instead complete graph, and (2) extracting only the essential parameters (6-dimensional quaternion pose and 1-dimensional scale) for each view within fragment, thereby avoiding costly global optimization of full point maps. Second, in the Hierarchical Gaussian Alignment (Global) stage, we merge fragments through a tree-based hierarchical framework that employs key frame guidance for inter-fragment alignment, pairwise merging with Gaussian point pruning, and subsequent joint optimization to ensure global consistency and mitigate cumulative errors efficiently.
@@
  • We introduce VideoLifter, an efficient, high-quality, and robust video-to-3D reconstruction framework with a local-to-global strategy.
  • Our fragment registration with learned 3D priors efficiently extracts essential representations for subsequent dense 3D-GS with several key efficiency-driven optimizations along with leverage learned priors to remove reliance on traditionally module SfM.
  • Our hierarchical 3D Gaussian alignment minimizes incremental errors through three well-designed iterative stages, ensuring both accuracy and efficiency.
  • Extensive experiments on the Tanks and Temples and CO3D-V2 datasets demonstrate that VideoLifter significantly enhances training efficiency, with more than 5× speed improvements, and improves rendering quality compared to state-of-the-art methods.
2 Related Works
@@
3D reconstruction for high-quality novel view synthesis is the task of generating unseen views of a scene or object from a set of given images [1, 29]. NeRF [30] is one of the most influential methods, but it remains computationally expensive and slow. Subsequent work has focused on either improving the quality of renderings [2, 3, 4] or boosting the efficiency of the approach [34, 37, 41, 40], though few methods succeed in both quality and efficiency simultaneously due to the ill-posed nature of transforming a sparse point cloud into a dense 3D scene representation. Recently, unstructured radiance field methods have emerged [7, 21, 48], using different scene representation primitives. Among these, 3D-GS [21] stands out and has demonstrated impressive performance in efficiently reconstructing complex, real-world scenes with high fidelity. Both NeRFs and 3DGS rely on carefully captured sequential video or multi-view images to ensure sufficient scene coverage, utilizing preprocessing tools like SfM software, e.g., COLMAP [36], to compute camera parameters and provide a sparse SfM point cloud as additional input. Specifically, 3D-GS performs the sparse-to-dense optimization starting from SfM point clouds, named adaptive density control, critical for converting sparse point clouds into dense 3D Gaussian representations.
@@
Estimating 3D structure and camera motion simultaneously is a well-explored challenge [15, 33, 47]. SfM has seen significant advancements across various dimensions. Methods like [26, 17] focus on enhancing feature detection, while [39] introduces innovative optimization approaches. Improved data representations and more robust structural solutions are explored by [16, 36]. Simultaneous Localization and Mapping (SLAM) systems, on the other hand, track camera motion and construct 3D structures from video sequences [33, 11, 32]. Despite these advances, traditional SfM and SLAM techniques remain vulnerable to issues such as low-texture regions, occlusions, moving objects, and lighting variations, limiting their overall robustness and performance, though some works [25] aim to improve it. Recent papers [43, 9, 10, 38] finds using NN-based methods to estimate matching/pose estimation, which finds more robust than the hand-designed feature SIFT.
@@
Inaccuracies from SfM can propagate through subsequent radiance field reconstruction, reducing overall quality. Various approaches have been proposed to eliminate the reliance on SfM by jointly optimizing camera parameters and scene representation. For instance, NeRFmm [46] and BARF [24] optimize camera intrinsics/extrinsics alongside the radiance field, while SCNeRF [19] address camera distortions via specialized parameterizations. GARF [19] further simplifies the joint optimization and improve both efficiency and accuracy by using Gaussian-MLP models. SPARF [44] and TrackNeRF [27] introduces a method to simulate pose noise by injecting Gaussian noise into the camera parameters.
Recently, depth priors from monocular depth estimator are used to guide radiance fields optimization [6, 8, 28, 14]. More recent work, such as [20] and InstantSplat [12], either supplement depth priors with other priors (e.g., image matching network), or integrate end-to-end stereo models like DUST3R/MASt3R to reduce the dependency on camera pose information. While these methods show promise in removing the SfM reliance, scaling them effectively to a large number of views remains a challenge.
Figure 2: CF-3DGS’s frame-by-frame pipeline accumulates errors, causing long-term drift, while our method could largely alleviate it. Results are tested on 247_26441_50907 from CO3D-V2.
Figure 3: Network Architecture. Given uncalibrated images, VideoLifter first employs learned priors for efficient fragment registration. The independently optimized 3D Gaussians from fragment are then hierarchically aligned into a globally coherent 3D representation.
@@ 3 VideoLifter: An Efficient and Effective Video-to-3D Framework
3.1 Video-to-3D: Challenges and Our Design
We first define the video-to-3D reconstruction problem and outline current issues in delivering an efficient and high-quality reconstruction pipeline. We then present our high-level design philosophy that tackles these challenges.
Video-to-3D Reconstruction:
Given a sequence of $N$ unposed and uncalibrated images from a monocular video, denoted as =$\mathcal{I}=\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$, VideoLifter aims to reconstruct the scene using 3D Gaussians $\mathcal{G}$ along with estimated camera intrinsics $K$ and extrinsics $\mathcal{T}=\{T_{i}\in\mathbb{R}^{3\times 4}\}_{i=1}^{N}$. We assume that all frames share a common intrinsic matrix, as they are from a single monocular video.
Key Challenges:
While NeRF and 3DGS have advanced 3D reconstruction, their variants remain suboptimal for video-to-3D reconstruction in terms of speed and quality. Existing methods often adopt a frame-by-frame progressive reconstruction paradigm, making them inherently slow and prone to cumulative errors when processing long sequence videos. As a result, long-term performance drift often occurs, as illustrated in Fig. 2. Furthermore, they typically rely on SfM to estimate camera intrinsic, which is unreliable or even infeasible for in-the-wild video sequences.
Our High-level Framework:
To meet the critical need for efficiency and quality in long-sequence video to 3D reconstruction, we depart from conventional frame-by-frame or holistic optimization methods by embracing a hierarchical local-to-global design philosophy. Specifically, we process long video sequences on a fragment basis and subsequently merge these fragments into a single, consistent 3D scene. Although the local-to-global concept has been explored in domains such as SLAM, adapting it to a video-to-3D reconstruction pipeline with 3DGS is new with two key, unanswered challenges:
  • How can we extract 3D reconstruction information efficiently and reliably from monocular video?
  • How can we merge multi-fragments into a high-quality and consistent 3D scene without alignment issues?
In response, we propose an efficient and high-quality video-to-3D reconstruction pipeline, VideoLifter, as illustrated in Fig.3. Our method achieves SOTA performance in both speed and quality (see Fig.1) through two key contributions: Local: Fragment Registration with Learned 3D Priors with efficiency-driven designs (Sec. 3.2) and Global: Hierarchical Gaussian Alignment to compress alignment error (Sec. 3.3).
@@ 3.2 Fragment Registration with Learned 3D Priors
Here, we describe our method to efficiently and reliably extract essential 3D reconstruction information on a fragment basis and the way we enforce cross-fragment consistency.
Process Description: We partition the input video sequence into $m$ disjoint windows of length $k$, which we refer to as fragments. For example, the $i$’th fragment is given by $\mathcal{I}^{f}_{i}=[I_{(i-1)k+1},I_{(i-1)k+2},\dots,I_{ik}]$, where $i\in[1,m]$. Our objective is twofold: first, to extract essential information for subsequent intra-fragment 3D reconstruction (e.g., point cloud for Gaussian initialization, along with coarse local camera extrinsics and intrinsics); and second, to obtain inter-fragment information necessary for future local-to-global merging.
Fragmentation Method Choice: In this work, we use a straightforward fragmentation strategy by uniformly dividing the frames into disjoint windows. Although simple, this approach has proven effective on benchmark datasets. More sophisticated methods, such as using frame-to-frame similarity to guide fragmentation, are orthogonal to our core contributions and could further enhance VideoLifter, e.g., in videos with abrupt view changes during capture.
Fragment-level Challenges: Naively applying existing methods for the fragment level (e.g., those in LocalRF [28] or CF-3DGS [14]) is neither efficient enough nor does it adequately prepare for future merging. Challenge ➀: SfM is heavily relied for NeRF/3D-GS, while it is not always available or reliable—especially in our video-to-3D reconstruction with varying conditions. Challenge ➁: A critical issue in the local-to-global paradigm is ensuring that fragments can be merged without incurring significant alignment errors. Challenge ➂: Beyond the inherent efficiency benefits of a local-to-global design, further enhancements in efficiency are necessary at the video setup.
Learned 3D priors (Challenge ➀): 3D-GS needs point cloud for initialization and camera pose for optimization. However, in long-sequence video settings, traditional SfM methods (e.g., COLMAP) are often unavailable or unreliable. While CF-3DGS employs monocular depth estimation (a geometric prior) on each view to obtain a point cloud, it introduces scale issues that necessitate additional optimization during 3D-GS training.
Inspired by recent work on replacing SIFT with NN-based method, e.g., LoFTR [43]/GIM [38], we abandon reliance on SfM and instead leverage learned 3D priors from large-scale pretrained foundation models. In this work, we use MASt3R [23] as prior model since it seamlessly integrates both geometric and matching cues. We emphasize that our video-to-3D pipeline are not exclusively tailored to MASt3R; rather, our VideoLifter is flexible and can incorporate any model that provides robust geometric and matching priors. For example, in Tab. 1, we demonstrate that VideoLifter performs well with both MASt3R and alternative approaches (e.g. LoFTR [43] for matching cues, Metric3Dv2 [18] for geometric cues). We choose MASt3R for its simplicity and efficiency.
Matching Geometric SSIM PSNR LPIPS ATE
LoFTR [43] Metric3Dv2 [18] 0.9238 31.30 0.0757 0.005
MASt3R MASt3R 0.9347 31.59 0.0730 0.004
Table 1: Comparison of prior models on Tanks and Temples.
Key Frames as Anchor (Challenge ➁, ➂): In a fragment-based approach, ensuring inter-fragment consistency is critical for high-quality 3D reconstruction. To address this, we propose to “anchor” each fragment with a key frame, set to its first frame, $I_{(i-1)k+1}$ and enforce consistency only among these key frames. This strategy simplifies the consistency problem by limiting the problem scale in $\frac{N}{k}$ key frames instead of considering whole $N$ frames, thereby improve efficiency and reduce complexity.
Key frame consistency is enforced at the point-cloud level by generating a globally optimized dense point map, along with transformation matrices across adjacent fragments $\{T^{f}_{i\rightarrow i+1}\}_{i=1}^{m-1}$. To achieve this, we follow a procedure similar to that in MASt3R [23], optimizing:
@@
where for each image pair $e=(v,u)\in\mathcal{E}$, $\sigma_{e}$ is scale factor, $\bm{P}_{v,e}$ and $\bm{O}_{v,e}$ is the pointmap and confidence map of $v$, respectively. However, MASt3R builds a complete graph for this optimization, resulting in a complexity of $\mathcal{O}((\frac{N}{k})^{2})$, it becomes inefficient enough for long video sequences.
To enhance efficiency, we propose to build a more efficient sub-graph that only builds edges between the key frames and their four closest neighboring frames. This design is motivated by the observation that neighboring segments share greater co-visibility; hence, edges between key frames with distant neighbors can be safely pruned. This sub-graph greatly reduces the optimization complexity to $\mathcal{O}(4\frac{N}{k})$, which scales linearly with #frames $N$, while still showing high end-to-end 3D reconstruction quality.
Efficient intra-Fragment Feature Registration (Challenge ➂): Finally, we aim to obtain an initial estimate of the local camera poses and pointmaps along with depth scale factors within each non-overlapping fragment, which can accelerate and boost the quality of the subsequent 3D Gaussian construction. A naive solution is to follow InstantSplat [12] to use MASt3R to solve the intra-fragment problem, but it requires optimizing millions of points and camera poses, making it inefficient.
To enhance efficiency, we leverage the pre-obtained key frame information and obtain the needed information only considering pairwise relationships between the key frame and all subsequent frames in the same fragment. In this way, we only need to solve for 6-dimensional camera poses (in quaternion format) and a 1-dimensional scale factor for each view, which has proven sufficient to maintain end-to-end reconstruction quality at high efficiency. Moreover, we believe that this simplified, non-sequential matching approach can reduce the incremental errors that are commonly encountered in sequential matching.
Take first fragment $\mathcal{I}^{f}_{1}=\{I_{1},\dots,I_{k}\}$ as an example.
@@
Camera pose: We refine the relative camera poses within the fragment using initial pairwise estimates. First, we identify the intersection of 2D correspondences between the key frame $I_{1}$ and each subsequent frame from index 2 to $k$. This process yields a consistent set of correspondences across all frames in the fragment. Using these intersected 2D correspondences, we retrieve the corresponding 3D positions from the key frame, which were previously optimized during key frame processing. These 3D-2D correspondences are then input to PnP-RANSAC [13], refining the camera poses to ensure alignment with consistent 3D points across views within the fragment. Only a 6-dim quaternion pose is optimized instead of directly optimizing pointmaps.